3- Find 2501 mod 17 (The remainder of th
Answer #1

y 59 (mania) = (1319,5)med 7) = (1258969 (05) (max 10) = {12reca ) (3x)(media) = (156 (med 17) . ( 32 mod.1))) = (62) (15) =